### v1

# Relative descriptors for quantum agents

Abstract
We use the Heisenberg picture of Everettian quantum mechanics to work out the relative descriptors for the Frauchiger-Renner thought experiment. While Everettian mechanics reveals no paradox within the thought experiment, the decoherence-free setup offers an instructive branching tree. Traditionally, branching structures derived from the Schrödinger picture suggest that foliations are always sharply defined. However, the Heisenberg picture demonstrates that the tree contains regions with inherently non-sharp foliations, leading to the conclusion that, in Everettian mechanics, quantum agents possess a blurry memory. This analysis helps resolve several misconceptions stemming from the Schrödinger picture.
keywords:
Heisenberg picture, Everettian quantum mechanics, thought experiments, branching.
1 Introduction
Everettian quantum theory is a realist interpretation of unitary quantum theory that eliminates the arbitrariness and anthropocentric biases of standard interpretations. Measurement outcomes are explained by the relative state construction, which is most commonly understood in the Schrödinger picture. The relative state construction, introduced by Everett almost 80 years ago Everett1957 , has significantly evolved over time. This evolution is summarised in Refs. Dewitt_Bryce_2016 ; saunders2010 ; Wallace2012 ; Brown2019 and has been further advanced by contributions from Deutsch Deutsch1985 ; Deutsch1999 ; Deutsch2000 ; Kuypers2021 .
A central concept in Everettian quantum theory is the branching tree, which represents the unitary evolution of the universe. Due to this branching structure, the theory has been popularly (and somewhat controversially among Everettians) known as the many-worlds interpretation of quantum mechanics. Despite the popularity of the branching tree concept, the precise rules for dividing a system into relative states have never been fully clarified. This ambiguity has led to a lack of consensus among Everettians regarding the branching process. While some argue that decoherence is necessary to define branches saunders2010 , others disagree Kuypers2021 . The most intriguing thought experiments occur in branching structures without decoherence, as full quantum coherence allows for the clearest comparison between different interpretations Nurgalieva2020 .
Recent progress by Kuypers and Deutsch Kuypers2021 involved formulating the relative state construction in the Heisenberg picture. Unlike the Schrödinger picture, the Heisenberg picture offers a precise description of when composite systems are foliated. Additionally, it provides a clear definition of what constitutes an Everettian universe, addressing a previously unresolved issue Vaidman2021 . In the Schrödinger picture, relative states have been interpreted as leading to parallel universes, which are seen as autonomously evolving components of the state vector. This interpretation has led to the misconception that a sharp branching structure can always be constructed. However, Kuypers and Deutsch demonstrated that, in general, a sharp branching structure does not exist. Insisting on foliating systems into parallel branches where it is not possible inevitably leads to paradoxes Gerhard2019 .
The Schrödinger picture has led to many misconceptions regarding the foliation process in Everettian quantum theory. In contrast, the Heisenberg picture clarifies that the relative state construction is an idealised approximation scheme. The clarity provided by Heisenberg observables (descriptors) simplifies their application to any well-defined system. However, there is a notable lack of worked-out examples in the current literature.
Here, we work out the relative descriptors for the Frauchiger-Renner thought experiment Frauchiger2018 ; Nurgalieva2020 ; Nurgalieva2022 , which continues to be a topic of debate in the context of various interpretations Lazarovici2019 ; Polychronakos2024 ; delRio2024 . Despite ongoing discussions, this thought experiment has invigorated the foundations of quantum mechanics, demonstrated the involvement of physicists in the realm of quantum interpretations, and spurred advancements in the field.
The experiment operates within a small Hilbert space and involves four quantum agents that communicate with each other. As we will show, because the protocol is an extended Wigner’s friend scenario, it serves as a prime example where a sharp branching tree is not feasible. This makes it an excellent exercise for working out the relative foliations for the thought experiment, while also addressing conceptual issues related to the foliation process in Everettian quantum theory.
The paper is organised as follows. In Sec. 2 we review state the thought experiment’s protocol in the form of a quantum circuit. In Sec. 3 we state the necessary elements from the Heisenberg picture of quantum computation relevant to the thought experiment. In Sec. 4 we work out the relative descriptors for the protocol, which allows us to construct the branching structure. We discuss the branching structure in Sec. 5.
2 The protocol
In this section, we outline the Frauchiger-Renner protocol. We assume the reader is familiar with the thought experiment, as it has been extensively discussed in the literature, not only by the original authors Frauchiger2018 ; Nurgalieva2020 ; Nurgalieva2022 , but also by others Lazarovici2019 ; DeBrota2020 ; Bub2021 ; Waaijer2021 .
The protocol may be represented by a quantum circuit; see Fig. 1. The system is modelled by eight qubits $\{R,A,S,B,U_{R},U_{A},W_{S},W_{B}\}$. Two labs are under the full quantum control of external agents. The first lab contains qubits $R$ and $A$, where $A$ represents Alice’s memory, and $R$ is a qubit in her lab. The second lab has qubits $S$ and $B$, where $B$ represents Bob’s memory, and $S$ is his qubit. Two external agents, Ursula $U$ and Wigner $W$ have full quantum control over Alice’s lab $RA$ and Bob’s lab $SB$, respectively. Ursula, to be able to measure both $R$ and $A$, must consist of at least two qubits herself. We call Ursula’s qubit that interacts with $R$ by $U_{R}$, and her other qubit that interacts with $A$ by $U_{A}$. Wigner has an analogous structure with $W_{S}$ and $W_{B}$. For brevity, we do not show Ursula’s and Wigner’s qubit legs in Fig. 1, and summarise their unitary interactions with the labs by measurement gates. A Qiskit implementation of the circuit in Fig. 1 may be found in Ref. Violaris_yt_23 .
Figure 1: The Frauchiger-Renner protocol.
We now describe the protocol of Fig. 1. All qubits initialise in the state |$\ket{0}$. In the Heisenberg picture, we may therefore represent the fixed Heisenberg state of the eight qubit network as |$\ket{\boldsymbol{0}}$. In the time interval $t\in(0,1)$, Alice’s qubit $R$ evolves through a rotation about the $y$ axis $R_{y}(\varphi)$ by and angle $\varphi=2\arcsin(\sqrt{2/3})$. Next, at $t\in(1,2)$, Alice’s unitary measurement on $R$ is modelled by a controlled-not gate, through which Alice records the eigenvalue $+1$ or $-1$ in her memory. Depending on the outcome, Alice prepares Bob’s qubit $S$. If Alice’s memory is $+1$, she leaves $S$ intact. But if Alice’s memory is $-1$, then she operates a Hadamard gate $H$ on $S$. In the circuit, this is achieved by a controlled-Hadamard gate at $t\in(2,3)$, with $A$ as the control, and $S$ as the target. At $t\in(3,4)$, Bob measures his qubit $S$ by means of a controlled-not gate. Between $t\in(4,7)$, Ursula and Wigner use the Bell-basis to measure Alice’s and Bob’s lab, respectively. The Bell-basis measurement is achieved by a controlled-not followed by a Hadamard Nielsen2010 . The implementation of this circuit shows that $U_{A}$ and $W_{B}$ always record the $+1$ eigenvalue, whereas $U_{R}$ and $W_{S}$ record both eigenvalues ±$\pm 1$ upon repetition of the protocol.
3 Descriptors
The objective of our work is to determine the Heisenberg observables, or descriptors, for all qubits at all times. In this section, we introduce the necessary descriptor preliminaries pertinent to the Frauchiger-Renner protocol. Except for the controlled-Hadamard, details of this section’s content may be consulted in Refs. Deutsch2000 ; Horsman2007 ; Horsman2007n ; Kuypers2021 ; Bedard2021abc ; Bedard2021 ; bedard2023 ; cruz2024 .
A generic qubit $Q$ is described by time-dependent Heisenberg observables, which are called descriptors and denoted by $\boldsymbol{q}_{Q}^{(t)}$. Qubit descriptors have three components, which at time $t=0$ coincide with the Schrödinger observables, and may be initialised as $\boldsymbol{q}_{Q}^{(0)}=\left(\sigma_{x},\sigma_{y},\sigma_{z}\right)_{Q}$, where $(\sigma_{x},\sigma_{y},\sigma_{z})$ are the Pauli matrices. Descriptor components $q_{Qi}^{(t)}$ ($i=x,y,z$) obey the Pauli algebra at all times, such that
$q_{Qi}^{(t)}q_{Qj}^{(t)}=\delta_{ij}I+i\sum_{k}\epsilon_{ijk}q_{Qk}^{(t)},\quad\left[\boldsymbol{q}_{Q}^{(t)},\boldsymbol{q}_{Q^{\prime}}^{(t)}\right]=0,\quad(Q\neq Q^{\prime}),$ (1)
where $I$ is the identity, and $\epsilon_{ijk}$ is the Levi-Civita symbol. The algebra allows us to omit one of the descriptor components for brevity, since that component is easily recovered from the other two.
To analyse the Frauchiger-Renner protocol in the Heisenberg picture, we must know how the rotation, the Hadamard, the controlled-not, and the controlled-Hadamard evolve the descriptor components. Given a generic quantum gate $G$, its effect on a descriptor component is given by the unitary Heisenberg evolution
$G:\quad q_{Qi}^{(t+1)}=U_{G}^{\dagger}\left[\boldsymbol{q}_{Q_{1}}^{(t)},\dots,\boldsymbol{q}_{Q_{N}}^{(t)}\right]q_{Qi}^{(t)}\,U_{G}\left[\boldsymbol{q}_{Q_{1}}^{(t)},\dots,\boldsymbol{q}_{Q_{N}}^{(t)}\right],$ (2)
where $\left\{\boldsymbol{q}_{Q_{1}}^{(t)},\dots,\boldsymbol{q}_{Q_{N}}^{(t)}\right\}$ is the set of all descriptors in an $N$ qubit computational network, and $U_{G}\left[\boldsymbol{q}_{Q_{1}}^{(t)},\dots,\boldsymbol{q}_{Q_{N}}^{(t)}\right]$ is called the functional representation of $G$.
We begin with the single qubit gates. For the rotation about the $y$ axis by an angle $\varphi$ on a qubit $Q$, the functional representation reads
$U_{R_{y}(\varphi)}\left[\boldsymbol{q}_{Q}^{(t)}\right]=I\cos\left(\frac{\varphi}{2}\right)-iq_{Qy}^{(t)}\sin\left(\frac{\varphi}{2}\right).$ (3)
Then, we may use Eq. (2) together with the algebra in Eq. (1) to work out the effect of the rotation on the descriptor components:
$R_{y}\left(\varphi\right):\quad\boldsymbol{q}_{Q}^{(t+1)}=\left(q_{Qx}^{(t)}\cos\varphi+q_{Qz}^{(t)}\sin\varphi,q_{Qy}^{(t)},q_{Qz}^{(t)}\cos\varphi-q_{Qx}^{(t)}\sin\varphi\right).$ (4)
Similarly, the functional representation for the Hadamard on $Q$ is
$U_{H}\left[\boldsymbol{q}_{Q}^{(t)}\right]=\frac{q_{Qx}^{(t)}+q_{Qz}^{(t)}}{\sqrt{2}},$ (5)
and the Hadamard evolves $Q$ as:
$H:\quad\boldsymbol{q}_{Q}^{(t+1)}=\left(q_{Qz}^{(t)},-q_{Qy}^{(t)},q_{Qx}^{(t)}\right).$ (6)
We now turn to the evolution of the controlled operations, which act on a control qubit $C$ and a target $T$. The controlled-not gate evolves those qubits as:
[…]
$\boldsymbol{q}_{T}^{(t+1)}$ $=\left(q_{Tx}^{(t)},q_{Ty}^{(t)}\underline{q_{Cz}^{(t)}},q_{Tz}^{(t)}\underline{q_{Cz}^{(t)}}\right),$ (7)
where the underlined terms highlight the parts that were copied from the other qubit during the interaction. The controlled-Hadamard has not been worked in the Heisenberg picture, but it is straightforward to do so. We obtain:
[…]
$\boldsymbol{q}_{C}^{(t+1)}=\left(q_{Cx}^{(t)}U_{H}\left[\boldsymbol{q}_{T}^{(t)}\right],q_{Cy}^{(t)}U_{H}\left[\boldsymbol{q}_{T}^{(t)}\right],q_{Cz}^{(t)}\right);$
$\boldsymbol{q}_{T}^{(t+1)}=\left(q_{Tx}^{(t)}P_{1}\left[q_{Cz}^{(t)}\right]+q_{Tz}^{(t)}P_{-1}\left[q_{Cz}^{(t)}\right],q_{Ty}^{(t)}q_{Cz}^{(t)},q_{Tz}^{(t)}P_{1}\left[q_{Cz}^{(t)}\right]+q_{Tx}^{(t)}P_{-1}\left[q_{Cz}^{(t)}\right]\right),$ (8)
where
$P_{\pm 1}\left[q_{Cz}^{(t)}\right]=\frac{1}{2}\left(I\pm q_{Cz}^{(t)}\right)$ (9)
are the projectors in the Heisenberg picture and form a projection-valued measure.
In the Heisenberg picture, a composite system may only be foliated into parallel universes if it is entangled. Entanglement is established through local interactions. Thus, what are known in the Schrödinger picture as parallel universes only spread through local interactions and have finite spatial extent. The Heisenberg picture shows that the foliations are better thought of as local bubbles that display Everettian multiplicity Kuypers2021 . Two qubits $Q$ and $Q^{\prime}$ are said to be entangled at time $t$, if there exists a pair of descriptor components such that Horsman2007 ; Horsman2007n ; Kuypers2021
$\left\langle q_{Qi}^{(t)}q_{Q^{\prime}j}^{(t)}\right\rangle\neq\left\langle q_{Qi}^{(t)}\right\rangle\left\langle q_{Q^{\prime}j}^{(t)}\right\rangle.$ (10)
[…]
A two-party composite system, such as a control and target qubit, may be sharply foliated into relative states at a time $t$ if it satisfies two conditions. First, the expectation values of the projectors are non-zero:
$\left\langle P_{\pm 1}\left[q_{Cz}^{(t)}\right]\right\rangle\neq 0\quad\mbox{and}\quad\left\langle P_{\pm 1}\left[q_{Tz}^{(t)}\right]\right\rangle\neq 0.$ (11)
Second, the product of the $z$ components of the control and target is sharp, that is
$\left\langle q_{Cz}^{(t)}q_{Tz}^{(t)}\right\rangle=1.$ (12)
With these conditions, one then may define foliated target $(T_{1},T_{-1})$ and control $(C_{1},C_{-1})$ qubit instances described by their corresponding relative descriptors
$T_{\pm 1}:\quad\boldsymbol{q}_{T,\pm 1}^{(t)}:=\boldsymbol{q}_{T}^{(t)}P_{\pm 1}\left[q_{Cz}^{(t)}\right];$
$C_{\pm 1}:\quad\boldsymbol{q}_{C,\pm 1}^{(t)}:=\boldsymbol{q}_{C}^{(t)}P_{\pm 1}\left[q_{Tz}^{(t)}\right],$ (13)
which define the foliations $C_{1}/T_{1}$ and $C_{-1}/T_{-1}$. For each foliation, one then defines a corresponding conditional expectation value. For instance, the conditional expectation value for the target, provided that the $z$ component of the control was measured to be ±$\pm 1$ is
$\left\langle q_{Ti}^{(t)}\right\rangle_{T,\pm 1}:=\frac{\left\langle q_{Ti,\pm 1}^{(t)}\right\rangle}{\left\langle P_{\pm 1}\left[q_{Cz}^{(t)}\right]\right\rangle}.$ (14)
The relative descriptors obey a Pauli algebra of reduced dimensionality.
4 Relative descriptors for the protocol
We now work out the descriptors for all times of the protocol illustrated in Fig. 1. For each time step, we then examine whether the emerging foliations are sharp or non-sharp. This will allow us to draw the Everettian branching tree, which unlike in the Schrödinger picture, will contain sharp and blurry parts. The algebra in Eq. (1) allows us to omit the $y$ component of the descriptors for brevity, which we may easily recover when necessary.
The qubit $R$ starts out described by $\boldsymbol{q}_{R}^{(0)}=(\sigma_{x},\sigma_{z})_{R}$, and so do all other seven qubits. In the time interval $t\in(0,1)$, the rotation $R_{y}(\varphi)$ evolves $R$ to
$\boldsymbol{q}_{R}^{(1)}=\left(\sigma_{x}c+\sigma_{z}s,\sigma_{z}c-\sigma_{x}s\right)_{R},$ (15)
where $c=-1/3$ and $s=\sqrt{8}/3$, which satisfy $c^{2}+s^{2}=1$. The expectation value $\left\langle q_{Rz}^{(1)}\right\rangle=c$ shows that $R$ lost its sharpness.
During $t\in(1,2)$, $A$ and $R$ interact through a controlled-not gate, with $R$ as the control, and $A$ as the target. Using Eq. (7), we find
$\boldsymbol{q}_{R}^{(2)}=\left[\left(\sigma_{x}^{R}c+\sigma_{z}^{R}s\right)\sigma_{x}^{A},\sigma_{z}^{R}c-\sigma_{x}^{R}s\right];$
$\boldsymbol{q}_{A}^{(2)}=\left[\sigma_{x}^{A},\sigma_{z}^{A}\left(\sigma_{z}^{R}c-\sigma_{x}^{R}s\right)\right].$ (16)
With Eq. (10) we may verify that $R$ and $A$ are now entangled. Moreover, given that the projectors have non-zero expectation value and that the product of the $z$ components is sharp, the composite system $RA$ at $t=2$ admits sharp foliations with relative descriptor components given by
$q_{Az,\pm 1}^{(2)}=q_{Az}^{(2)}P_{\pm 1}\left[q_{Rz}^{(2)}\right]=\pm\sigma_{z}^{A}P_{\pm 1}\left[q_{Rz}^{(2)}\right].$ (17)
It is sufficient to write the relative descriptor of the $z$ component of the target. Using Eq. (14) we see that the conditional expectation values $\left\langle q_{Az}^{(2)}\right\rangle_{A,\pm 1}=\pm 1$ are sharp. The analysis determines the two sharp foliations $R_{\pm 1}/A_{\pm 1}$.
Based on these sharp outcomes, $A$ prepares the qubit $S$. The interaction between them is a controlled-Hadamard between $t\in(2,3)$, with $S$ as the target. Using Eq. (8), we find
$\boldsymbol{q}_{A}^{(3)}=\left(\sigma_{x}^{A}U_{H}\left[\boldsymbol{q}_{S}^{(2)}\right],q_{Az}^{(2)}\right);$
$\boldsymbol{q}_{S}^{(3)}=\left(\sigma_{x}^{S}P_{1}\left[q_{Az}^{(2)}\right]+\sigma_{z}^{S}P_{-1}\left[q_{Az}^{(2)}\right],\sigma_{z}^{S}P_{1}\left[q_{Az}^{(2)}\right]+\sigma_{x}^{S}P_{-1}\left[q_{Az}^{(2)}\right]\right).$ (18)
We may work out the descriptor components in terms of the initial Pauli matrices. However, because the algebra of the descriptors is valid at all times, it is not always advantageous nor necessary to do so. $A$ and $S$ are entangled at $t=3$, because
$\left\langle q_{Az}^{(3)}q_{Sz}^{(3)}\right\rangle\neq\left\langle q_{Az}^{(3)}\right\rangle\left\langle q_{Sz}^{(3)}\right\rangle.$ (19)
Nonetheless, the product $\left\langle q_{Az}^{(3)}q_{Sz}^{(3)}\right\rangle=1/3$, which does not yield sharp foliations. Not all entangled systems foliate parallelly, which is hard to see in the Schrödinger picture. Note that at $t=3$, $R_{\pm 1}/A_{\pm 1}$ remains sharp. It is the $A/S$ part that forms an interference bubble.
At $t\in(3,4)$, $B$ measures his qubit $S$ through a controlled-not operation, with $B$ as the target. The descriptors at time $t=4$ are
$\boldsymbol{q}_{S}^{(4)}=\left(q_{Sx}^{(3)}\sigma_{x}^{B},q_{Sz}^{(3)}\right);\qquad\boldsymbol{q}_{B}^{(4)}=\left(\sigma_{x}^{B},\sigma_{z}^{B}q_{Sz}^{(3)}\right).$ (20)
This allows for the relative instances
$q_{Bz,\pm 1}^{(4)}=\pm\sigma_{z}^{B}P_{\pm 1}\left[q_{Sz}^{(4)}\right],$ (21)
defining the sharp foliations $S_{\pm 1}/B_{\pm 1}$.
During $t\in(4,6)$, Ursula and Wigner perform a Bell-measurement on the $RA$ and $SB$ labs, respectively. The first stage includes a controlled-not between $R$ and $A$, and between $S$ and $B$, followed by a Hadamard on $R$ and $S$. This evolves the involved qubits to
$\boldsymbol{q}_{R}^{(6)}=\left(q_{Rz}^{(2)},q_{Rx}^{(2)}q_{Ax}^{(3)}\right);\quad\boldsymbol{q}_{A}^{(6)}=\left(q_{Ax}^{(3)},q_{Az}^{(2)}q_{Rz}^{(2)}\right);$
$\boldsymbol{q}_{S}^{(6)}=\left(q_{Sz}^{(3)},q_{Sx}^{(4)}\sigma_{x}^{B}\right);\quad\boldsymbol{q}_{B}^{(6)}=\left(\sigma_{x}^{B},q_{Bz}^{(4)}q_{Sz}^{(3)}\right).$ (22)
This stage (before $U$ and $W$ interact) diffuses the $R_{\pm 1}/A_{\pm 1}$ and $S_{\pm 1}/B_{\pm 1}$ foliations that immediately before were sharp.
In the last step at $t\in(6,7)$, Ursula and Wigner interact with their respective labs. These interactions are controlled-not gates between $R$ and $U_{R}$, $A$ and $U_{A}$, $S$ and $W_{S}$, and $B$ and $W_{B}$. In Fig. 1, these interactions are condensed into the measurement gates. The resulting descriptors for all eight qubits are
$\boldsymbol{q}_{R}^{(7)}=\left(q_{Rz}^{(2)}\sigma_{x}^{U_{R}},q_{Rx}^{(2)}q_{Ax}^{(3)}\right);\quad\boldsymbol{q}_{U_{R}}^{(7)}=\left(\sigma_{x}^{U_{R}},\sigma_{z}^{U_{R}}q_{Rx}^{(2)}q_{Ax}^{(3)}\right);$
$\boldsymbol{q}_{A}^{(7)}=\left(q_{Ax}^{(3)}\sigma_{x}^{U_{A}},q_{Az}^{(2)}q_{Rz}^{(2)}\right);\quad\boldsymbol{q}_{U_{A}}^{(7)}=\left(\sigma_{x}^{U_{A}},\sigma_{z}^{U_{A}}q_{Az}^{(2)}q_{Rz}^{(2)}\right);$
$\boldsymbol{q}_{S}^{(7)}=\left(q_{Sz}^{(3)}\sigma_{x}^{W_{S}},q_{Sx}^{(4)}\sigma_{x}^{B}\right);\quad\boldsymbol{q}_{W_{S}}^{(7)}=\left(\sigma_{x}^{W_{S}},\sigma_{z}^{W_{S}}q_{Sx}^{(4)}\sigma_{x}^{B}\right);$
$\boldsymbol{q}_{B}^{(7)}=\left(\sigma_{x}^{B}\sigma_{x}^{W_{B}},q_{Bz}^{(4)}q_{Sz}^{(3)}\right);\quad\boldsymbol{q}_{W_{B}}^{(7)}=\left(\sigma_{x}^{W_{B}},\sigma_{z}^{W_{B}}q_{Bz}^{(4)}q_{Sz}^{(3)}\right).$ (23)
All qubits at $t=7$ admit sharp relative descriptors, which for the final targets are
$q_{U_{R},z,\pm 1}^{(7)}=\pm\sigma_{z}^{U_{R}}P_{\pm 1}\left[q_{Rz}^{(7)}\right];\quad q_{U_{A},z,1}^{(7)}=\sigma_{z}^{U_{A}}P_{1}\left[q_{Az}^{(7)}\right];$
$q_{W_{S},z,\pm 1}^{(7)}=\pm\sigma_{z}^{W_{S}}P_{\pm 1}\left[q_{Sz}^{(7)}\right];\quad q_{W_{B},z,1}^{(7)}=\sigma_{z}^{W_{B}}P_{1}\left[q_{Bz}^{(7)}\right].$ (24)
Note that $U_{A}$ and $W_{B}$ are guaranteed to observe the $+1$ eigenvalue. $U_{R}$ and $W_{B}$ may distinguish between Bell states. The resulting foliations are $R_{\pm 1}/{U_{R}}_{\pm 1}/A_{1}/{U_{A}}_{1}$ and $S_{\pm 1}/{W_{S}}_{\pm 1}/B_{1}/{W_{B}}_{1}$.
We summarise this section’s results in Tab. 1. In the last column we also state all the expectation values for the projectors for all interactions at each time step. These results are in agreement with the projections known from the Schrödinger picture Nurgalieva2020 .
Time Interaction Gate Foliations Projections
$(0,1)$ Rotation on $R$
$(1,2)$ $R,A$ Controlled-not Sharp $1/3,2/3$
$(2,3)$ $A,S$ Controlled-$H$ Non-sharp $1/3,2/3$
$(3,4)$ $S,B$ Controlled-not Sharp $1/3,2/3$
$(4,5)$ $R,A$ Controlled-not Non-sharp $1/3,2/3$
$(4,5)$ $S,B$ Controlled-not Non-sharp $1/3,2/3$
$(5,6)$ $R,A$ Hadamard on $R$ Non-sharp $5/6,1/6$
$(5,6)$ $S,B$ Hadamard on $S$ Non-sharp $5/6,1/6$
$(6,7)$ $R,U_{R}$ Controlled-not Sharp $5/6,1/6$
$(6,7)$ $A,U_{A}$ Controlled-not Sharp $1,0$
$(6,7)$ $S,W_{S}$ Controlled-not Sharp $5/6,1/6$
$(6,7)$ $B,W_{B}$ Controlled-not Sharp $1,0$
Table 1: Summary of the time intervals $(t_{1},t_{2})$, the parties involved $(X,Y)$, the operation on the parties, the sharpness of the resulting foliations, and the values of the expectation values of the projections $\left\langle P_{1}\left[q_{Xz}^{(t_{2})}\right]\right\rangle$ and $\left\langle P_{-1}\left[q_{Xz}^{(t_{2})}\right]\right\rangle$.
5 Discussion
Based on Tab. 1, in Fig. 2 we draw the Everettian branching tree for the Frauchiger-Renner protocol. In contrast to the Schrödinger picture, here the branching tree has non-sharp parts and all foliations are spatially local. In the interval $t\in(0,1)$ the Hadamard changes the expectation value of the $R$ qubit, which is indicated by a colour change. The black trunk of the tree contains instances of all eight qubits, but we emphasise the specific subsystems that are being foliated. If $R$ and $A$ foliate, all other systems remain unaffected. The first sharp foliation occurs at $t\in(1,2)$, when the controlled-not originates $R_{\pm 1}/A_{\pm 1}$. The projection weights of these foliations may be consulted in Tab. 1. These foliations remain unaffected until $t\in(4,5)$, when another controlled-not at the same physical location diffuses $R_{\pm 1}/A_{\pm 1}$. Between $t\in(2,3)$, the controlled-Hadamard among $A$ and $S$ leads to the first interference bubble, which we indicate by the blurry part with the colour reflecting the projection weights from the sharp foliations entering and emerging from the region. Whatever enters the blurry region, contributes to the outcomes of foliations that emerge from it at a later time. Note that until $t\in(3,4)$, the same instance of $S$ and $B$ were carried through all branches of the tree until they finally foliated into $S_{\pm 1}/B_{\pm 1}$ due to a controlled-not at their physical location. At $t\in(4,5)$ both the $R_{\pm 1}/A_{\pm 1}$ and $S_{\pm 1}/B_{\pm 1}$ foliations diffuse through controlled-not gates. Between $t\in(5,6)$ the Hadamard on $R$ and $S$ changes the projections, which we again indicate with the colour transition. The final measurements by Ursula and Wigner lead to the final foliations $R_{\pm 1}/{U_{R}}_{\pm 1}/A_{1}/{U_{A}}_{1}$ and $S_{\pm 1}/{W_{S}}_{\pm 1}/B_{1}/{W_{B}}_{1}$.
Figure 2: Everettian branching tree based on the relative descriptor construction for the Frauchiger-Renner protocol.
The Frauchiger-Renner though experiment stirred the communities’ interest, because they showed that at least one of three seemingly conservative statements must be false. We recall the statements here: (Q) Any agent may describe any system around them using the formalism of quantum theory and, based on this description, infer predictions (or retrodictions) of measurement outcomes using the quantum-mechanical Born rule; (C) An agent can use conclusions obtained by admitting the view of another agent; (S) A measurement carried out by an agent has a definitive outcome from the viewpoint of the measuring agent. It is known that Everettian quantum theory rejects at least one of the rules Frauchiger2018 . In light of the tree in Fig. 2, we may comment on the statements in the present context. Here, (S) remains true. To evaluate (C), the original argument uses Alice’s relative outcome $R_{-1}/A_{-1}$ to infer that $B_{1}/{W_{B}}_{1}$. Yet, quantum theory admits both $B_{\pm 1}/{W_{B}}_{\pm 1}$. The reason is that both foliations $R_{\pm 1}/A_{\pm 1}$ contribute to $B_{\pm 1}/{W_{B}}_{\pm 1}$ through a non-sharp region. The same happens for the pilot-wave in Bohmian mechanics Lazarovici2019 . The contradiction with quantum mechanics only arises of one insists having sharp foliations at all times, which is not possible in the Everettian theory. Wigner cannot adopt the viewpoint (C) of Alice/Bob, because they are using incompatible bases. Condition (Q) is emergent, but not fundamental in unitary quantum theory. The situation is nicely summarised by Ref’s. vedral2020local abstract: "Unperformed measurements have no results. Unobserved results can affect future measurements."
6 Conclusion
We worked out the relative descriptors for the Frauchiger-Renner thought experiment. In the context of Everettian quantum mechanics, the exercise shows that in experiments where agents have full quantum control over other agents, the branching tree will have both sharp and non-sharp parts. Sharp relative foliations may be defined from entangled composite systems, but not every entangled system foliates sharply. The conclusion one may draw from this is that quantum agents have a blurry memory.
[…]
Acknowledgements
The authors acknowledge the funding from CNPq.